THE LENGTH OF A RECTANGLE IS 14 FEET MORE THAN THE WIDTH. THE PERIMETER OF THE RECTANGLE IS 136 FIND THE LENGTH.
Answer by ohdearitslydia(1)

Length- x+14
Width- x
Formula for perimeter- 2x+2(x+14)=136
Distribute- 2x+2x+28=136
Add and subtract like terms- 4x=108
Divide- x=27
Plug it back in- x+14
27+14=41
Length=41